What concentration of the lead ion, Pb2+, must be exceeded to precipitate PbF2 from a solution that is 1.00×10−2M in the fluoride ion, F−? Ksp for lead(II) fluoride is 3.3×10−8 .
Answer #1

Q > K

Q >  3.3×10−8 .

Q = [Pb+2][F-]^2

if F- = 10^-2

Solve for Pb+2

Q = [Pb+2][F-]^2

3.3×10−8 . = (Pb+2) ( 10^-2 .)^2

[Pb+2] = 3.3*10^-4

Anything more of that concnetration will make a precipitate